OpenStudy (anonymous):

If x=y+5, Isn't The x And y Intercept -5.?

OpenStudy (saifoo.khan):

x is +5. while y is -5.

OpenStudy (anonymous):

How Is x Positive 5.?

OpenStudy (saifoo.khan):

When you have to find the "x" intercept, make y = 0 x = y + 5 x = 0 + 5 x = 5.

OpenStudy (anonymous):

Okay.Thanks.